Booster
- running well
- today back to 12 turns
- would like to go up to 5.5e12-6e12 per spill

BNB
- DI bottle starting to fail. On Wednesday swap the bottle, turn off beam 3am
4 hr cool off. probably back by noon. 
- time with two bottles didn't get any longer. 
- move raw skid upstream during one of the future shutdowns.

- commission trim magnet ~about half a shift and will be done before validating horn 4.

SciBooNE area
- got in touch with Josh, MITPC not taking data will remove detector from the hall
- ANNIE using neutron source Wednesday through today.
yesterday added PMT on top neutron capture vessel

MicroBooNE
- smooth run
- plan to run NuMI open data trigger when NuMI ramps up
- NuMI: running with recylcer starting this weekend. next week ramp up intensity